OpenStudy (mathmath333):

the mean of 1,2,2^2,2^3.......2^31 lies in between a. 2^24 to 2^25 b. 2^25 to 2^26 c. 2^26 to 2^27 d. 2^29 to 2^30 e. 2^23 to 2^24

OpenStudy (amistre64):

well, we have a geometric sum of 32 data points

OpenStudy (amistre64):

\[\frac{\left(\frac{1-r^{31}}{1-r}\right)}{32}\]

OpenStudy (mathmath333):

how comes this formula??

OpenStudy (amistre64):

a mean is the sum of the parts, divided by the number of parts: 2^0 to 2^31 is 32 parts the sum of the parts is geometric by definition: let S = r^0 + r^1 + r^2 + ... + r^31 and -rS = - r^1 - r^2 - ... - r^31 - r^32 -------------------------------------- (1-r)S = r^0 - .................................... - r^32 so S = (1 - r^32)/(1-r) i spose i had a slight error due to the last term

OpenStudy (amistre64):

r^n, given n terms .... and n=32 in this case :)

OpenStudy (amistre64):

at any rate: \[\frac{1-2^{32}}{2^5(1-2)}=2^{27}-2^{-5}\]

OpenStudy (amistre64):

looks like 2^27 - a fraction, so its between ^26 and ^27

OpenStudy (mathmath333):